### 24 Jan 2024 v2

@@ 1 Introduction
Social media platforms, while holding the potential to facilitate communication and foster informed discussions, are also susceptible to the dissemination of misinformation and disinformation campaigns [6, 7]. This issue extends beyond politics and seeps into sensitive domains like public health, as exemplified by the anti-vaccine movements during the COVID-19 pandemic [8]. Compelling evidence abounds, pointing to the active exploitation of social media platforms by certain governments to subvert domestic social movements and interfere in the democratic elections of foreign adversaries [9]. Noteworthy instances of such foreign interventions include the case of the 2017 French presidential election [10] and the highly significant interference by the Internet Research Agency (IRA: a Russian company engaged in online influence operations on behalf of Russian business and political interests) in the 2016 US presidential election [11, 12].
@@
Our findings provide evidence supporting this notion. A consistent group of suspended accounts exhibits similarity with IRA accounts in terms of the information they interact with and disseminate to the broader Twitter community. We demonstrate that this group of suspended accounts did influence (in a causal sense) the retweet activity of undecided users and weak supporters—individuals uncertain about their voting choices—in terms of political polarization.
The paper is structured as follows: In Section 2.1, we provide information about the data utilized and the methods employed. Section 3 delves into the results of our analyses, following these steps: a) Characterizing users in our dataset based on the information they disseminate; b) Employing the IRA ego network as a proxy to identify a subset of suspended users exhibiting similar patterns; and c) Assessing the impact of this group of suspended accounts on Twitter discourse through Granger causality. The paper concludes with a comprehensive discussion and conclusion that encapsulates insights derived from our analyses.
@@ 2.3 Retweet network
A link between two users occurs every time a user $u$ retweets a user’s tweet $v$ that contains a URL linking to a website belonging to one of the news media categories. The direction of the connection goes from $v$ to $u$, i.e., the direction of the information flow between Twitter users. We do not include multiple links in the same direction between the same two users, nor do we include self-links. The degree of a node in the network is defined as the number of edges connected to it. The out-degree of a node $u$, $k_{out}^{u}$, represents the number of unique users who retweeted $u$. On the other hand, the in-degree of a node $u$, $k_{in}^{u}$, represents the number of users retweeted by node $u$. It is worth noticing that, by construction, these networks are balanced directed networks, and as such, $\langle k_{in}\rangle=\langle k_{out}\rangle=\langle k\rangle/2$.
@@ 2.5 Two-sample Kolmogorov-Smirnov test
Version one-sided To test whether non-IRA users are more likely to interact with IRA users or vice-versa, we perform a two-sample Kolmogorov-Smirnov test between the two distributions extracted by each interaction network. The null hypothesis, denoted as $H_{0}$, assumes that the activity of a given user in each interaction type, represented by $x=(x_{out},x_{in})$, follows the condition $F_{out}(x)\geq F_{in}(x)$ for every $x$. Here, $F_{out}(x)$ and $F_{in}(x)$ represent the cumulative density functions (CDF) for the “out” and “in” directions, respectively. The alternative hypothesis, on the other hand, suggests that $F_{out}(x)<F_{in}(x)$ for at least one $x$.
@@
Version two-sided Similarly to the one-sided version, to test for differences in the in/out-degree activity of suspended, not found, not verified, verified, and IRA accounts, we employed a Two-sample Kolmogorov-Smirnov test with null hypothesis $H_{0}$: $F_{i}(x)=F_{j}(x)$ where $x=k_{type}^{s}$. The superscript $s$ indicates that the degree considered comes from the sampled nodes, $i,j\in$ (suspended, not found, verified, not verified, IRA), and $type\in$ (in, out). We set a level of 5%, meaning that we will reject the null hypothesis and favor the alternative if the p-value is less than 0.05.
@@ 3.1 Accounts characterization
Finally, we test whether suspended, not found, verified, not verified, and IRA behave differently in terms of their in/out activity in each news category network(see Section 2.3 and Table 1), we employ a two-sample Kolmogorov-Smirnov test [27, 28] (two-sided version, see Section 2.5) with null hypothesis $H_{0}$: the data are drawn from the same distribution. We performed the test for each two-pair combination of the groups $(\overrightarrow{k_{type}}^{i},\overrightarrow{k_{type}}^{j})$, with $i,j\in$ (suspended, not found, verified, not verified, IRA), and $type\in$ (in, out). The $\overrightarrow{k_{type}}$ vector contains the values of $\langle k_{type}\rangle$ for each category network. We adapt sampling strategies to avoid sampling bias error, as explained in Section 2.4. Refer to Supplementary Table 7 for a view of the sample degrees.
(a)Refer to caption
(b)Refer to caption
Figure 3: Two-sample Kolmogorov-Smirnov tests. Results of the two sample Kolmogorov Smirnov test between groups based on their out degree (a) and in degree (b) are presented. Yellow boxes indicate rejection of the null hypothesis in favor of the two-sided alternative, suggesting that the data were not drawn from the same distribution. In panel (a), we observe that $H_{0}$ is only rejected between IRA and suspended accounts, indicating a similarity in the out-degree activity of these two groups. Furthermore, verified accounts exhibit an out-degree activity that differs from that of the other groups. Analyzing the in degree activity, we find that verified accounts only behave similarly to IRA accounts.
Figures 3a and b show the results of the tests for the out-degree and in-degree, respectively. We used a heatmap representation, where the yellow color indicates the rejection of the null hypothesis $H_{0}$ in favor of the default two-sided alternative, suggesting that the data were not drawn from the same distribution.
@@ 3.2 IRA ego network
To pinpoint a representative subset of suspended accounts employing strategies akin to IRA accounts, we delve into the IRA ego network. This network encompasses all users interacting with IRA through retweets, mentions, quotes, or replies (see Section 2.3). Table 2 summarizes the key features of these networks.
@@
Figure 4: Account status: top 1000 active users. We show the fraction of the top 1000 active users per status (in %) in each interaction network (retweet in blue, mention in red, reply in gree, and quote in gray). Possible statuses are: Verified, Suspended, Not Verified and Not Found. (a) “out” direction: users are retweeted, mentioned, quoted, or replied to by IRA users. On average, 18.4% of these users have verified accounts, 49.4% are not verified, 18.9% are suspended, and 11.4% are not found accounts (b) “in” direction: users retweet, mention, quote, or reply to IRA users. On average, 45.1% of them are not verified, 35.7% is suspended, and 17.1% is not found.
@@
Ego polarization Analyzing user preferences in terms of political orientations (see Sections 2.6 and 2.7) reveals, as expected, a higher presence of right supporters in the IRA ego network. Specifically, 8.1% of users are strong supporters of Trump, 4% are strong supporters of Clinton, 64.5% are weak supporters of Trump, 21.1% are weak supporters of Clinton, and the remaining 2.3% of users are categorized as undecided.
The majority of suspended accounts fall into the category of Trump supporters, whether weak or strong, as shown in Fig. 3. This aligns with the notion that this subset of suspended accounts was oriented toward the right agenda. Additional details about the users’ classifications can be found in Supplementary Section 4.
@@
The distinction in the political orientation of these two communities is further supported by analyses of the hashtag clouds constructed from the content shared by users, as shown in Fig. 5. Users in the right community tend to share hashtags in support of Trump and against Clinton, while the opposite holds true for the left community. See Supplementary Section 6.
(a)Refer to caption
(b)Refer to caption
Figure 5: Hashtags clouds. (a) Hashtags clouds for community right. The prevalence of hashtags such as #neverHillary, #maga, #trump, and #benghazi suggests that the community has a strong affinity toward Donald Trump. (b) Hashtags cloud for community left. Prominent hashtags are #nevertrump, #notmypresident, #imwithher, and #hillaryclinton, suggesting strong affinity toward Hillary Clinton. Additionally, this word cloud includes hashtags associated with the black community, such as #blacklivesmatter and #blacktwitter, referring to ideologies in opposition to the ideologies espoused by Trump.
Within the right community, suspended accounts comprise 20% of the total nodes in the community. This percentage decreases by more than half in the left community (refer to Table 4). This outcome suggests a difference in strategies between suspended accounts and IRA, with IRA implementing a dual strategy (targeting both right and left users), while suspended accounts predominantly focus on targeting right-oriented users. These suspended users account for 21.7% of the interactions, with the directions of the connections going from IRA to suspended accounts, as indicated in Table 5 and Supplementary Figure 6. In the case of the left community, the type of interaction is inverted, and in most cases, it is the IRA engaging with the other groups, particularly not verified and verified, with the connections going from suspended to IRA accounting for 12.5% of the overall connections in the community.
@@ 3.3 Causal network patterns: IRA nodes versus suspended nodes
This section delves into measuring the impact that suspended accounts and IRA have on shaping the Twitter discourse during the 2016 US presidential elections. Specifically, we scrutinize the causal relationships between IRA (and suspended accounts) tweet activity and the activity of supporting classes, namely weak Trump supporters, weak Clinton supporters, strong Trump supporters, strong Clinton supporters, and undecided users.
@@
Tables 8 and 9 present the causal relationships among different groups in the two scenarios: one with only IRA nodes and the other with suspended nodes. The direction of each link is from the column group to the row group. For example, considering the strong Trump supporters, their causal effect on the weak Clinton supporters is measured at 0.16 $\pm$ 0.011, as shown in Table 8. The blue entries in the tables represent the auto-correlation of each time series. In both scenarios, the auto-correlations exhibit the strongest causal effects for all time series, except for the undecided group.
To identify the most significant causal links, a threshold of 0.16 (0.20 for the suspended scenario) was set on the causal relation, selecting connections that account for 75% of the total effect. These selected links are highlighted in bold in the tables. Figures 6a and 6b visualize the causal networks constructed using these connections. The nodes are colored as follows: dark red for strong Trump supporters, dark blue for strong Clinton supporters, light red for weak Trump supporters, light blue for weak Clinton supporters, orange for the IRA nodes, and gray for the undecided group. Arrows indicate the direction of maximal causal effect ($\geq$0.16 and $\geq$ 0.20) between two activity time series. The width of each arrow represents the strength of the causation, and the size of each node is proportional to the auto-correlation of each time series.
@@
Figure 6: Causal Networks. (a) Graph showing the maximal causal effects between the activity of the IRA nodes and the supporting classes of the presidential candidates. Arrows indicate the direction of the maximal causal effect ($\geq 0.16$) between two activity time series. The width of each arrow is proportional to the strength of the causation, and the size of each node is proportional to the auto-correlation of each time series. Dark blue and dark red highlight the contribution of strong Clinton and Trump supporters, respectively. Light blue and light red are associated with the weak Clinton and Trump supporters, gray with the undecided users, and orange with the IRA nodes. The causal relation primarily flows from strong supporters of both Trump and Clinton to weak and strong supporters of opposing political candidates. Additionally, weak supporters from both sides play a role in influencing the undecided group, with weak Trump supporters receiving support from strong Trump supporters in their efforts. Notably, IRA nodes do not play a significant role in this causal network, suggesting that they have limited influence on shaping Twitter discourse. (b) Graph showing the maximal causal effects between the activity of the suspended nodes and the supporting classes of the presidential candidates. Arrows indicate the direction of the maximal causal effect ($\geq 0.2$) between two activity time series. Strong Trump supporters have a causal effect on IRA+S nodes, which, in turn, have a causal influence on both weak supporters and the undecided group. Additionally, weak supporters continue to exert a causal effect on the undecided group. Strong Trump supporters have a causal effect on strong Clinton supporters, but not vice-versa.
@@ 4 Discussion
Having identified a group of suspended accounts resembling IRA behaviors (partially) and gaining insights into their intent, we proceed to measure the impact of both suspended accounts and IRA on shaping the Twitter discourse during the 2016 US presidential elections. This is achieved through the application of Granger causality to the tweet activity produced by IRA suspended accounts, and each supporting class (refer to Section 3.3).
@@
It’s important to note that while Granger causality suggests that past retweet behaviors in one group can aid in predicting the retweet behaviors of the other group, it doesn’t imply a direct cause-and-effect relationship. The determination of such causality goes beyond the scope of this study. Additionally, utilizing tweet activity provides insights into user behavior, but conclusions regarding changes in users’ vote intentions require longitudinal data.
@@
In summary, this study suggests a scenario in which a significant group of suspended accounts, often overshadowed by the IRA narrative, played a crucial role during the 2016 US presidential elections. Further research is required to delve deeper into their impact on political user preferences.
@@
IRA, Internet Research Agency;
@@
HAM was supported by NSF-HNDS Award 2214217. JSA gratefully acknowledges the Brazilian agencies FUNCAP, CNPq and CAPES, the National Institute of Science and Technology for Complex Systems in Brazil, and the PRONEX-FUNCAP/CNPq Award PR2-0101-00050.01.00/15 for financial support. Matteo Serafino gratefully acknowledges the Brazilian agency CAPES Award 88887.899221/2023-00.
@@
HAM, AB and MS conceived the research; MS and AB designed and supervised the research; MS and AB coordinated and supervised the analysis; MS and ZZ performed the analyses. MS and ZZ, JSA, AB, and HAM analyzed the results; MS and ZZ wrote the first draft. All authors edited and approved the paper.

### 13 Mar 2024 v3

@@ 1 Introduction
Social media platforms, while holding the potential to facilitate communication and foster informed discussions, are also susceptible to the dissemination of misinformation and disinformation campaigns [6, 7, 8]. This issue extends beyond politics and seeps into sensitive domains like public health, as exemplified by the anti-vaccine movements during the COVID-19 pandemic [9]. Compelling evidence abounds, pointing to the active exploitation of social media platforms by certain governments to subvert domestic social movements and interfere in the democratic elections of foreign adversaries [10]. Noteworthy instances of such foreign interventions include the case of the 2017 French presidential election [11] and the highly significant interference by the Internet Research Agency (IRA: a Russian company engaged in online influence operations on behalf of Russian business and political interests) in the 2016 US presidential election [12, 13].
@@
Our findings provide evidence supporting this notion. A consistent group of suspended accounts exhibits similarity with IRA accounts in terms of the information they interact with and disseminate to the broader Twitter community. We demonstrate that the group of suspended accounts did indeed influence, in a Granger-causal manner, the retweet activity of undecided users and weak supporters—individuals uncertain about their voting decisions—in terms of political polarization.
The paper is organized as follows: Section 2 outlines the data collection and analysis methods utilized throughout the study. In Section 3, we present the findings of our research, which include: a) the characterization of users within our dataset based on the content they share; b) the use of the IRA ego network as a means to identify a group of suspended users with similar behavioral patterns; and c) an evaluation of how this specific group of suspended accounts influences Twitter discourse, utilizing Granger causality for assessment. The manuscript concludes with a thorough discussion and conclusion section, summarizing the key insights gained from our analysis.
@@ 2.3 Retweet network
In the context of a news category network, a link between two users occurs every time a user $u$ retweets a user’s tweet $v$ that contains a URL linking to a website belonging to one of the news media categories. The direction of the connection goes from $v$ to $u$, i.e., the direction of the information flow between Twitter users. We do not include multiple links in the same direction between the same two users, nor do we include self-links. The degree of a node in the network is defined as the number of edges connected to it. The out-degree of a node $u$, $k_{out}^{u}$, represents the number of unique users who retweeted $u$. On the other hand, the in-degree of a node $u$, $k_{in}^{u}$, represents the number of users retweeted by node $u$. It is worth noticing that, by construction, these networks are balanced directed networks, and as such, $\langle k_{in}\rangle=\langle k_{out}\rangle=\langle k\rangle/2$.
@@ 2.5 Two-sample Kolmogorov-Smirnov test
To test for differences in the in/out-degree activity of suspended, not found, not verified, verified, and IRA accounts, we employed a Two-sample Kolmogorov-Smirnov test with null hypothesis $H_{0}$: $F_{i}(x)=F_{j}(x)$ where $x=k_{type}^{s}$. The superscript $s$ indicates that the degree considered comes from the sampled nodes, $i,j\in$ (suspended, not found, verified, not verified, IRA), and $type\in$ (in, out). The null hypothesis, denoted as $H_{0}$, assumes that the activity of a given user in each interaction type, represented by $x=(x_{out},x_{in})$, follows the condition $F_{out}(x)=F_{in}(x)$ for every $x$. Here, $F_{out}(x)$ and $F_{in}(x)$ represent the cumulative density functions (CDF) for the “out” and “in” directions, respectively. The alternative hypothesis, on the other hand, suggests that $F_{out}(x)<F_{in}(x)$ (or $F_{out}(x)>F_{in}(x)$) for at least one $x$.
@@ 3.1 Accounts characterization
Finally, we test whether suspended, not found, verified, not verified, and IRA behave differently in terms of their in/out activity in each news category network (see Section 2.3 and Table 1). We employ a two-sample Kolmogorov-Smirnov test [28, 29] (two-sided version, see Section 2.5) with null hypothesis $H_{0}$: the data are drawn from the same distribution. We performed the test for each two-pair combination of the groups $(\overrightarrow{k_{type}}^{i},\overrightarrow{k_{type}}^{j})$, with $i,j\in$ (suspended, not found, verified, not verified, IRA), and $type\in$ (in, out). The $\overrightarrow{k_{type}}$ vector contains the values of $\langle k_{type}\rangle$ for each category network. We adapt sampling strategies to avoid sampling bias error, as explained in Section 2.4. Refer to Supplementary Table 7 for a view of the sample degrees. While the news category networks considered here are not weighted, we did check for differences in the weighted case and could not find any significant variations.
Figure 3: Two-sample Kolmogorov-Smirnov tests. Results of the two sample Kolmogorov Smirnov test between groups based on their out degree. Yellow boxes indicate rejection of the null hypothesis in favor of the two-sided alternative, suggesting that the data were not drawn from the same distribution. We observe that $H_{0}$ is only rejected between IRA and suspended accounts, indicating a similarity in the out-degree activity of these two groups. Furthermore, verified accounts exhibit an out-degree activity that differs from that of the other groups.
Figure 3 shows the results of the tests for the out-degree and in-degree, respectively. We used a heatmap representation, where the yellow color indicates the rejection of the null hypothesis $H_{0}$ in favor of the default two-sided alternative, suggesting that the data were not drawn from the same distribution.
@@ 3.2 IRA ego network
To pinpoint a representative subset of suspended accounts employing strategies akin to IRA accounts, we look into the IRA ego network. This network encompasses all users interacting with IRA through retweets, mentions, quotes, or replies (see Section 2.3). Table 2 summarizes the key features of these networks.
@@
Figure 4: Account status: top 1000 active users. We show the fraction of the top 1000 active users per status (in %) in each interaction network (retweet in blue, mention in red, reply in green, and quote in gray). Possible statuses are: Verified, Suspended, Not Verified and Not Found. (a) “out” direction: users are retweeted, mentioned, quoted, or replied to by IRA users. On average, 18.4% of these users have verified accounts, 49.4% are not verified, 18.9% are suspended, and 11.4% are not found accounts (b) “in” direction: users retweet, mention, quote, or reply to IRA users. On average, 45.1% of them are not verified, 35.7% is suspended, and 17.1% is not found.
@@
Ego polarization Analyzing user preferences in terms of political orientations (see Sections 2.6 and 2.7) reveals, as expected, a higher presence of right supporters in the IRA ego network. Specifically, 8.1% of users are strong supporters of Trump, 4% are strong supporters of Clinton, 64.5% are weak supporters of Trump, 21.1% are weak supporters of Clinton, and the remaining 2.3% of users are categorized as undecided. It’s important to note that when calculating user polarization in computing, we consider all users, regardless of their account status. The majority of suspended accounts are classified as Trump supporters, either weak or strong. This is detailed in Table 3, where we also provide the percentages broken down by account status.
This aligns with the notion that this subset of suspended accounts was oriented toward the right agenda. Additional details about the users’ classifications can be found in Supplementary Section 4.
@@
The distinction in the political orientation of these two communities is further supported by analyses of the hashtag clouds constructed from the content shared by users, as discussed in Supplementary Section 6. Users in the right community tend to share hashtags in support of Trump and against Clinton, while the opposite holds true for the left community.
Within the right community, suspended accounts comprise 20% of the total nodes in the community. This percentage decreases by more than half in the left community (refer to Table 4). This outcome suggests a difference in strategies between suspended accounts and IRA, with IRA implementing a dual strategy (targeting both right and left users), while suspended accounts predominantly focus on targeting right-oriented users. These suspended users account for 21.7% of the interactions, with the directions of the connections going from IRA to suspended accounts, as indicated in Supplementary Figure 7 and Supplementary Table 20. In the case of the left community, the type of interaction is inverted, and in most cases, it is the IRA engaging with the other groups, particularly not verified and verified, with the connections going from suspended to IRA accounting for 12.5% of the overall connections in the community.
@@ 3.3 Causal network patterns: IRA nodes versus suspended nodes
This section investigates the impact that suspended accounts and IRA have on shaping the Twitter discourse during the 2016 US presidential elections. Specifically, we scrutinize the causal relationships between IRA (and suspended accounts) tweet activity and the activity of supporting classes, namely weak Trump supporters, weak Clinton supporters, strong Trump supporters, strong Clinton supporters, and undecided users.
@@
Tables 6 and 7 present the causal relationships among different groups in the two scenarios: one with only IRA nodes and the other with suspended nodes. The direction of each link is from the column group to the row group. For example, considering the strong Trump supporters, their causal effect on the weak Clinton supporters is measured at 0.16 $\pm$ 0.011, as shown in Table 6. The blue entries in the tables represent the auto-correlation of each time series. In both scenarios, the auto-correlations exhibit the strongest causal effects for all time series, except for the undecided group.
To identify the most significant causal links, a threshold of 0.16 (0.20 for the suspended scenario) was set on the causal relation, selecting connections that account for 75% of the total effect. These selected links are highlighted in bold in the tables. Figures 5a and 5b visualize the causal networks constructed using these connections. The nodes are colored as follows: dark red for strong Trump supporters, dark blue for strong Clinton supporters, light red for weak Trump supporters, light blue for weak Clinton supporters, orange for the IRA nodes, and gray for the undecided group. Arrows indicate the direction of maximal causal effect ($\geq$0.16 and $\geq$ 0.20) between two activity time series. The width of each arrow represents the strength of the causation, and the size of each node is proportional to the auto-correlation of each time series.
@@
Figure 5: Causal Networks. (a) Graph showing the maximal causal effects between the activity of the IRA nodes and the supporting classes of the presidential candidates. Arrows indicate the direction of the maximal causal effect ($\geq 0.16$) between two activity time series. The width of each arrow is proportional to the strength of the causation, and the size of each node is proportional to the auto-correlation of each time series. Dark blue and dark red highlight the contribution of strong Clinton and Trump supporters, respectively. Light blue and light red are associated with the weak Clinton and Trump supporters, gray with the undecided users, and orange with the IRA nodes. The causal relation primarily flows from strong supporters of both Trump and Clinton to weak and strong supporters of opposing political candidates. Additionally, weak supporters from both sides play a role in influencing the undecided group, with weak Trump supporters receiving support from strong Trump supporters in their efforts. Notably, IRA nodes do not play a significant role in this causal network, suggesting that they have limited influence on shaping Twitter discourse. (b) Graph showing the maximal causal effects between the activity of the suspended nodes and the supporting classes of the presidential candidates. Arrows indicate the direction of the maximal causal effect ($\geq 0.2$) between two activity time series. Strong Trump supporters have a causal effect on Suspended nodes, which, in turn, have a causal influence on both weak supporters and the undecided group. Additionally, weak supporters continue to exert a causal effect on the undecided group. Strong Trump supporters have a causal effect on strong Clinton supporters, but not vice-versa.
@@ 4 Discussions and Conclusions
Having identified a group of suspended accounts resembling IRA behaviors (partially) and gaining insights into their intent, we proceed to measure the impact of both suspended accounts and IRA on shaping the Twitter discourse during the 2016 US presidential elections. This is achieved through the application of Granger causality to the tweet activity produced by IRA and suspended accounts, and each supporting class (refer to Section 3.3).
@@
It’s important to note that while Granger causality suggests that past retweet behaviors in one group can aid in predicting the retweet behaviors of the other group, it doesn’t imply a direct cause-and-effect relationship. The determination of such causality goes beyond the scope of this study. Additionally, utilizing tweet activity provides insights into user behavior, but conclusions regarding changes in users’ vote intentions require longitudinal data. It is, however, noteworthy that the similarity observed between the Internet Research Agency (IRA) and the group of suspended users, coupled with the fact that suspended accounts influenced the activity of undecided users, opens up the possibility of a new scenario, such as potential cooperation between the IRA and the identified group of suspended users. It’s also conceivable that this group of suspended accounts was part of the IRA’s campaign and remained undetected by Twitter. However, this remains purely speculative, and further analysis and data are needed to draw more concrete conclusions.
@@
In summary, this study suggests a scenario in which a significant group of suspended accounts, often overshadowed by the IRA narrative, played a crucial role during the 2016 US presidential elections. Further research is required to better understand their impact on political user preferences.
@@
IRA: refer to the Internet Research Agency.
@@
HAM was supported by NSF-HNDS Award 2214217. JSA gratefully acknowledges the Brazilian agencies FUNCAP, CNPq and CAPES, the National Institute of Science and Technology for Complex Systems in Brazil, and the PRONEX-FUNCAP/CNPq Award PR2-0101-00050.01.00/15 for financial support. MS gratefully acknowledges the Brazilian agency CAPES Award 88887.899221/2023-00. ZZ was supported by the National Natural Science Foundation of China under project No. 62302319 and R&D Program of Beijing Municipal Education Commission (Grant No. KM202210038002).
@@
HAM, AB and MS conceived the research; MS and AB designed and supervised the research; MS and AB coordinated and supervised the analysis; MS and ZZ performed the analyses. MS, ZZ, JSA, AB, and HAM analyzed the results; MS and ZZ wrote the first draft. All authors edited and approved the paper.
Acknowledgements
Not applicable.